OpenStudy (anonymous):

Identify the vertex for the graph of y = 2x2 + 8x − 3.

OpenStudy (anonymous):

The overall formula to find the x-coordinate of the vertex, based on y=ax^2+bx+c, is -b/2a. All you have to do is plug in the a and the b to their respective spots: -8/4 which simplifies to -2. Now that is your x coordinate of the vertex. All you have to do is plug x into the equation and solve for y: y= 2(-2)^2 + 8(-2) -3 y= 8 - 16 - 3 y= -11 Your vertex is (-2, -11)

OpenStudy (anonymous):

wow, thanks flixoe